Aresearch institute poll asked respondents if they felt vulnerable to identity theft. In the poll, n=1064 and x=572 who said yes.Use a 99 % confidence level . Find the best point estimate of the population proportion p
Answer by Boreal(15235)

The point estimate mss 572/1064=0.5376
the half-interval for a CI=99% is z(0.995)*sqrt(0.5376*0.4624/1064)
=2.576*0.0153=0.0393
so the interval is (0.4983, 0.5769)
Where one rounds will affect the fourth decimal place.